The length of time needed to complete a certain test is normally distributed with mean 23 minutes and standard deviation 11 minutes. Find the probability that it will take more than 20 minutes to complete the test.
Answer #1

Solution :

Given that ,

mean = = 23

standard deviation = = 11

P(x > 20) = 1 - p( x< 20 )

=1- p P[(x - ) / < (20 - 23) / 11 ]

=1- P(z < - 0.27)

Using z table,

= 1 - 0.3936

= 0.6064